OpenStudy (anonymous):

The struggle is real! Find the time t in years when the annual sales x of a new product are increasing at the greatest rate. Use a graphing utility to verify your results.

OpenStudy (dan815):

whats the function

OpenStudy (anonymous):

it keeps kicking me out of the problem =/

OpenStudy (anonymous):

\[\frac{ 900000t^2 }{ 144+t^2 }\]

OpenStudy (anonymous):

@dan815

OpenStudy (anonymous):

So far I have (144+ t^2)(1800000t)-(900000t^2)(2t)

OpenStudy (anonymous):

hello!

OpenStudy (anonymous):

When they say graphing utility, do they want you to plug it into the graphing calculator to find the largest slope?

OpenStudy (anonymous):

because "increasing at the greatest rate" part

OpenStudy (anonymous):

I think so, but I get error when I plug it in ,because its such a big number

OpenStudy (anonymous):

hmm do you have a ti81?

OpenStudy (anonymous):

ti-84 plus

OpenStudy (anonymous):

oh :/ you might just plug in manually with mathematics and then just drawing it out and estimate

OpenStudy (aum):

Factor 10^5 out and keep it as a constant outside.

OpenStudy (anonymous):

you just need values I think to justify that it's increasing at the greatest rate so, maybe. But yeah @aum that seems to be the best choice.

OpenStudy (aum):

If you want to solve this mathematically first and later verify your answer with the graphing calculator, then: The first derivative gives the rate of increase. For it to be a maximum, take the second derivative, equate it to zero and solve for t.

OpenStudy (anonymous):

I'm going to try that next I have the problem I believe to to the equate zero part and solve for t and got 3sqrt2, but webassign marked it wrong .

OpenStudy (anonymous):

err I got to the equation zero part I think I did probably something wrong at 259200000(144+t^2)^2-(4t^3+576t)(259200000t)

OpenStudy (anonymous):

I know for sure though the bottom is (144+t^2)^3

OpenStudy (aum):

\[ f(t) = \frac{ 900000t^2 }{ 144+t^2 } = 10^5 * \frac{ 9t^2 }{ 144+t^2 } \\ f'(t) = 10^5 * \frac{ (144+t^2)*18t-9t^2(2t) }{ (144+t^2)^2 } = 10^5 * \frac{ 2592t + 18t^3 - 18t^3 }{ (144+t^2)^2 } = \\ f'(t) = 10^5 * \frac{ 2592t }{ (144+t^2)^2 } \\ \]

OpenStudy (aum):

\[ f'(t) = 10^5 * \frac{ 2592t }{ (144+t^2)^2 } \\ f''(t) = 10^5 * \frac{(144+t^2)^2(2592) - 2592t(2)(144+t^2)(2t)}{(144+t^2)^4} = 0\\ \]

OpenStudy (aum):

\[ 2592(144+t^2)(144+t^2 - 4t^2) = 0 \\ 2592(144+t^2)(144- 3t^2) = 0 \\ 144 = 3t^2 \\ t^2 = 48 = 16 * 3 \\ t = 4\sqrt{3} \\ \]

OpenStudy (anonymous):

WOW thank you so much!

OpenStudy (aum):

You are welcome.

OpenStudy (anonymous):

I have been struggling with this problem for the last two hours

OpenStudy (aum):

Glad to be able to help.

OpenStudy (anonymous):

Took a screen shot for future use:) Thanks again:)